# Modeling fast X-ray variability around an accreting black hole

### v1

@@ 5.1 The roles of the propagating fluctuation and the light traveling in the inter-band lag
In Sect. 4, Fig. 4 shows a low-frequency hard lag and a high-frequency soft lag between 0.1-0.3 keV where disk, Compton, and reflection components are comparable, and 2-5 keV where there are only continuum emissions of Compton and reflection components. We briefly analyze the viscously propagating lag, hereafter propagation lag, caused by the delay in propagating fluctuation, and the spatial lag caused by the delay in light travel. Here, we describe how the spatial lags and propagation lags result in the lag spectra with Fig. 7.
@@
Figure 7: The corona obstructs the partial accretion disc, and due to the emission profile on the disc, the disc component originates from the outer region while the Comptonized photons come from the inner region. This results in fluctuations of the Compton and reflection components occurring on the timescale of the inner region. The propagation lag between the outer and inner regions, which occurs over several seconds, is greater than the light-travel delay that occurs over milliseconds. As a result, the propagation lag predominates in the low-frequency band, leading to the observed low-frequency hard lag. Additionally, this propagation lag diminishes at higher frequencies, where spatial lags take precedence. The travel time for soft thermal photons reaching the corona is several dozen $R_{\rm g}/c$, which is shorter than the travel time for hard Comptonized photons interacting with the disc, measured in hundreds of $R_{\rm g}/c$. Consequently, at high frequencies, this longer travel delay for hard photons results in the soft lag.
@@ 5.2 Observational test: explanation of the evolution of the lags in MAXI J1820+070
The observed frequency-dependent lag shows a trend that the zero-crossing points shift toward higher frequency and the amplitudes of soft lags decrease, indicating the evolution of the inner accretion flow (Kara et al., 2019; Marco et al., 2021). These observational results can be successfully reproduced in the simulations in this work, by changing both the geometry of the corona and the maximum viscous frequency (Fig. 5). In this section, we will show that the evolution of the lag spectrum can be attributed to the change of both the geometry and the maximum viscous frequency, as is shown in Fig. (6).
The changes in geometry play a role in the evolution of the lag at high frequency. As demonstrated in Fig. 6(a), the zero-crossing point shifts toward higher frequencies as the height of the corona decreases. As the separation between the corona and the disc decreases (see Fig. 8(a)), the spatial lag for the Comptonized photons striking the disc also diminishes, resulting in a reduction of the soft lag. Furthermore, a shorter soft lag causes its zero-crossing point to shift toward higher frequencies (Mizumoto et al., 2018; Marco et al., 2021; Yu et al., 2023). Thus, the decrease in spatial lag due to the geometrical change of the corona accounts for the reduced amplitude of the lag and the shift of the zero-crossing points to higher frequencies, as illustrated in Fig. 5.
@@ 5.3 Alternative spherical corona case
Further observations and simulations are needed to distinguish the explicit geometry of the corona clearly. X-ray polarization detection is a highly effective method for probing this geometry. Recent findings from the Imaging X-ray Polarimetry Explorer (IXPE) indicate that the corona is extended perpendicular to the resolved radio jet for several X-ray binaries (Krawczynski et al., 2022; Ingram et al., 2024). The spherical corona model, which can explain the timing results as well as any other geometry, is thus ruled out by IXPE results, because it predicts the coronal emission to be unpolarized. The jet base model, in contrast, can be consistent with IXPE results if the parameters are chosen to ensure the corona is wider than it is tall, as they are this paper. The two mechanisms we explore in this paper to vary the soft lag amplitude will give rise to very different polarization signals. The polarization should stay constant as the maximum viscous frequency increases, whereas the polarization degree is expected to increase as the coronal height reduces, thus making the corona even more vertically extended. Dense IXPE monitoring of future X-ray binary outbursts will therefore provide a powerful diagnostic. IXPE monitoring of Swift J1727.8-1613 demonstrated that the aspect ratio of the corona remains almost constant during the hard intermediate state, even though the soft lag amplitude increases dramatically (Ingram et al., 2024). However, the epoch of the hard state in which the soft lag amplitude reduces was already over before the first IXPE observation took place. With the upcoming launch of the Enhanced X-ray Timing and Polarimetry mission (eXTP), additional polarimetry observations will become available, necessitating further simulations of X-ray polarization to refine our understanding of the corona’s geometry. Therefore, we plan to conduct polarimetry simulations for various extended corona models in our future research.
@@ 6 Acknowledgment
We thank Stefano Rapisarda for his help in modeling the propagating fluctuation, and Phil Uttley for his suggestion regarding the implication of the simulation. We also thank Wei Yu, Liang Zhang, and Wenda Zhang for their helpful discussion/comments. We also thank Yongkang Zhang for his help in preparing the schematic. B.Y. is supported by Natural Science Foundation of China (NSFC) grants 12322307, 12361131579, and 12273026; by the National Program on Key Research and Development Project 2021YFA0718500; by Xiaomi Foundation / Xiaomi Young Talents Program. FYW is supported by the Natural Science Foundation of China (NSFC) grants 12494575 and 12273009. A.I. acknowledges support from the Royal Society.

### v2

@@ 5.1 The roles of the propagating fluctuation and the light traveling in the inter-band lag
In Sect. 4, Fig. 4 shows . We briefly analyze the viscously propagating lag, hereafter propagation lag, caused by the delay in propagating fluctuation, and the spatial lag caused by the delay in light travel. Here, we describe how the spatial lags and propagation lags result in the lag spectra with Fig. 7.
@@
Figure 7: The colors on the disc serve as a visual representation of the variability speed across the disc, where a deeper yellow intensity indicates a higher viscous frequency. The corona obstructs the partial accretion disc, and due to the emission profile on the disc, the disc component originates from the outer region, while the Comptonized photons come from the inner region. This results in fluctuations of the Compton and reflection components occurring on the timescale of the inner region. The propagation lag between the outer and inner regions, which occurs over several seconds, is greater than the light-travel delay that occurs over milliseconds. As a result, the propagation lag predominates in the low-frequency band, leading to the observed low-frequency hard lag. Additionally, this propagation lag diminishes at higher frequencies, where spatial lags take precedence. The travel time for soft thermal photons reaching the corona is several dozen $R_{\rm g}/c$, which is shorter than the travel time for hard Comptonized photons interacting with the disc, measured in hundreds of $R_{\rm g}/c$. Consequently, at high frequencies, this longer travel delay for hard photons results in the soft lag.
@@ 5.2 Observational test: explanation of the evolution of the lags in MAXI J1820+070
The observed frequency-dependent lag shows a trend that the zero-crossing points shift toward higher frequency and the amplitudes of soft lags decrease, indicating the evolution of the inner accretion flow (Kara et al., 2019; De Marco et al., 2021). These observational results can be successfully reproduced in the simulations in this work, by changing both the geometry of the corona and the maximum viscous frequency (Fig. 5). In this section, we will show that the evolution of the lag spectrum can be attributed to the change of both the geometry and the maximum viscous frequency, as is shown in Fig. (6).
The changes in geometry play a role in the evolution of the lag at high frequency. As demonstrated in Fig. 6(a), the zero-crossing point shifts toward higher frequencies as the height of the corona decreases. As the separation between the corona and the disc decreases (see Fig. 8(a)), the spatial lag for the Comptonized photons striking the disc also diminishes, resulting in a reduction of the soft lag. Furthermore, a shorter soft lag causes its zero-crossing point to shift toward higher frequencies (Mizumoto et al., 2018; De Marco et al., 2021; Yu et al., 2023). Thus, the decrease in spatial lag due to the geometrical change of the corona accounts for the reduced amplitude of the lag and the shift of the zero-crossing points to higher frequencies, as illustrated in Fig. 5.
@@ 5.3 Alternative spherical corona case
Further observations and simulations are needed to distinguish the explicit geometry of the corona clearly. X-ray polarization detection is a highly effective method for probing this geometry. Recent findings from the Imaging X-ray Polarimetry Explorer (IXPE) indicate that the corona is extended perpendicular to the resolved radio jet for several X-ray binaries (Krawczynski et al., 2022; Ingram et al., 2024). The spherical corona model, which can explain the timing results as well as any other geometry, is thus ruled out by IXPE results, because it predicts the coronal emission to be unpolarized. The jet base model, in contrast, can be consistent with IXPE results if the parameters are chosen to ensure the corona is wider than it is tall, as they are in this paper. The two mechanisms we explore in this paper to vary the soft lag amplitude will give rise to very different polarization signals. The polarization should stay constant as the maximum viscous frequency increases, whereas the polarization degree is expected to increase as the coronal height reduces, thus making the corona even more vertically extended. Dense IXPE monitoring of future X-ray binary outbursts will therefore provide a powerful diagnostic. IXPE monitoring of Swift J1727.8-1613 demonstrated that the aspect ratio of the corona remains almost constant during the hard intermediate state, even though the soft lag amplitude increases dramatically (Ingram et al., 2024). However, the epoch of the hard state in which the soft lag amplitude reduces was already over before the first IXPE observation took place. With the upcoming launch of the Enhanced X-ray Timing and Polarimetry mission (eXTP), additional polarimetry observations will become available, necessitating further simulations of X-ray polarization to refine our understanding of the corona’s geometry. Therefore, we plan to conduct polarimetry simulations for various extended corona models in our future research.
@@ 6 Acknowledgment
We thank Stefano Rapisarda for his help in modeling the propagating fluctuation, and Phil Uttley for his suggestion regarding the implication of the simulation. We also thank Wei Yu, Liang Zhang, Wenda Zhang, and Honghui Liu for their helpful discussion/comments. We also thank Yongkang Zhang for his help in preparing the schematic. B.Y. is supported by Natural Science Foundation of China (NSFC) grants 12322307, 12361131579, and 12273026; by the National Program on Key Research and Development Project 2021YFA0718500; by Xiaomi Foundation / Xiaomi Young Talents Program. FYW is supported by the Natural Science Foundation of China (NSFC) grants 12494575 and 12273009. A.I. acknowledges support from the Royal Society.